Suppose a bank offers a savings account with the interest rate of 6% compounded annually. If you deposit $100, how long does it take to triple your money if you do not withdraw any from your account?
Answer #1

Future value = 100 * 3 = 300

Future value = Present value ( 1 + r)n

300 = 100 ( 1 + 0.06)n

3 = (1.06)n

LN 3 = n LN 1.06

1.098612 = n 0.058269

n = 18.85

It will take 18.85 years.